# FUIA: Model Inversion Attack against Federated Unlearning

### v2

@@ VIII-A Gradient Pruning
During local training on each client, we prune the least significant gradients to obscure gradient information, thereby reducing the effectiveness of FUIA. Specifically, we prune a proportion of the gradients with the smallest absolute values in each training round. As shown in Figure 13, when , the reconstructed images by FUIA become increasingly blurry as increases, with MSE values rising and PSNR values dropping. When , FUIA fails to reconstruct any critical information about the forgotten data, demonstrating the effectiveness of this method. However, as increases, the validation accuracy on the test set decreases, as shown in Table II, with accuracies of 76, 70, 62, and 53, respectively. To balance defense effectiveness and model performance, we select as the pruning ratio, which effectively defends against FUIA while controlling the decline in model performance.
@@ VIII-B Gradient Perturbation
Inspired by differential privacy, we add Gaussian noise to gradients during client training to obscure the target gradients, reducing the effectiveness of FUIA. Specifically, we add Gaussian noise with a mean of 0 and standard deviations () of . As shown in Figure 14, as the noise intensity ( value) increases, the reconstructed images by FUIA gradually lose their target features and become increasingly blurred, with MSE values rising and PSNR values dropping. When the value exceeds 0.003, the reconstructed images by FUIA lose most of the critical features of the forgotten data. However, adding noise also affects model performance, as shown in Table II, where validation accuracies under different noise intensities are 80, 71, 64, and 54, respectively. Therefore, we choose a Gaussian noise with as the defense parameter, which effectively defends against FUIA while preserving as much model performance as possible.
@@ IX Conclusion
In this study, we firstly focus on the privacy leakage issue inherent in FU and propose FUIA. FUIA is applicable to most existing FU methods and adapts its attack strategy based on different unlearning targets. The core mechanism involves tracking parameter changes during the unlearning process to reconstruct forgotten data. This attack directly contradicts the fundamental goal of FU, which is to completely remove the influence of specific data, thereby exposing inherent privacy vulnerabilities in FU methods. To evaluate the effectiveness of FUIA, we conduct extensive experiments demonstrating its ability to infer private information from forgotten data. Additionally, we perform a systematic ablation study to analyze the impact of key variables on attack performance. Furthermore, we explore two potential defense mechanisms to mitigate the security risks posed by FUIA. Our work highlight the privacy risks in FU methods and aim to contribute to the development of more secure and robust federated unlearning mechanisms.

### v3

@@ VIII-A Gradient Pruning
During local training on each client, we prune the least significant gradients to obscure gradient information, thereby reducing the effectiveness of FUIA. Specifically, we prune a proportion of the gradients with the smallest absolute values in each training round. As shown in Figure 14, when , the reconstructed images by FUIA become increasingly blurry as increases, with MSE values rising and PSNR values dropping. When , FUIA fails to reconstruct any critical information about the forgotten data, demonstrating the effectiveness of this method. However, as increases, the validation accuracy on the test set decreases, as shown in Table II, with accuracies of 78.32, 72.27, 67.78, and 62.21, respectively. To balance defense effectiveness and model performance, we select as the pruning ratio, which effectively defends against FUIA while controlling the decline in model performance.
@@ VIII-B Gradient Perturbation
Inspired by differential privacy, we add Gaussian noise to gradients during client training to obscure the target gradients, reducing the effectiveness of FUIA. Specifically, we add Gaussian noise with a mean of 0 and standard deviations () of . As shown in Figure 15, as the noise intensity ( value) increases, the reconstructed images by FUIA gradually lose their target features and become increasingly blurred, with MSE values rising and PSNR values dropping. When the value exceeds 0.003, the reconstructed images by FUIA lose most of the critical features of the forgotten data. However, adding noise also affects model performance, as shown in Table II, where validation accuracies under different noise intensities are 80.64, 76.43, 72.33, and 65.32, respectively. Therefore, we choose a Gaussian noise with as the defense parameter, which effectively defends against FUIA while preserving as much model performance as possible.
@@ IX Conclusion
In this study, we firstly focus on the privacy leakage issue inherent in FU and propose FUIA. FUIA covers the main current FU scenarios and adapts its attack strategy based on different unlearning targets. The core mechanism involves tracking parameter changes during the unlearning process to reconstruct forgotten data. This attack directly contradicts the fundamental goal of FU, which is to completely remove the influence of specific data, thereby exposing inherent privacy vulnerabilities in FU methods. To evaluate the effectiveness of FUIA, we conduct extensive experiments demonstrating its ability to infer private information from forgotten data. Additionally, we perform a systematic ablation study to analyze the impact of key variables on attack performance. Furthermore, we explore two potential defense mechanisms to mitigate the security risks posed by FUIA. Our work highlight the privacy risks in FU methods and aim to contribute to the development of more secure and robust federated unlearning mechanisms.